4y + square root 3 x = 8
3y + 4x = -6
Is there a solution - how do I graph this system? Please help.
Thanks

Answer by stanbon(75887)

1st: 4y + square root 3 x = 8
2nd: 3y + 4x = -6
---------------
Multiply 1st by 3 and 2nd by 4 to get:
3rd: 12y+3sqrt3 x = 24
4th: 12y+16x = -24
---------------
Subtract 4th from 3rd to get:
5th: (3sqrt3 -16)x=48
x=48/(3sqrt3 - 16)
x=-4.44286...
---------------
Substitute in 2nd to solve for y, as follows:
3y +4(-4.442861643)=-6
3y - 17.7714467...=-6
3y=11.77144657...
y=3.923815524...
------------
These value check in 3rd
------------------------
You should be able to graph the original equations as:
y=(1/4)(8-(sqrt3)x)
y=(1/3)(-6-4x)

Cheers,
Stan H.